The Federal Reserve System publishes data on family income based on its Survey of Consumer Finances. When the head of the household has a college degree, the mean before-tax family income is $ 84,950. Suppose that 58% of the before-tax family incomes when the head of the household has a college degree are between $74,800 and $95,100 and that these incomes are normally distributed. What is the standard deviation of before-tax family incomes when the head of the household has a college degree? (Round the value of z to 2 decimal places. Round your answer to 2 decimal places.)
Answer #1

\\ $ Mean $ (\mu) = \$84,950 \\\\ $ Let standard deviation be $ \sigma \\\\ $ 58\% are between \$74,800 and \$95,100 $ \\\\ \Rightarrow P(\ 74,800 < x < 95,100\ ) = 0.58 \\\\ $ As $ z = \dfrac{ x - \mu}{\sigma} \\\\ \Rightarrow P(\ \dfrac{74,800 - 84,950}{\sigma} < z < \dfrac{95,100 - 84,950}{\sigma}\ ) = 0.58 \\\\ \Rightarrow P(\ -\dfrac{10,150}{\sigma} < z < \dfrac{10,150}{\sigma}\ ) = 0.79 - 0.21 \\\\ \Rightarrow (\ -\dfrac{10,150}{\sigma} <z < \dfrac{10,150}{\sigma}\ )= P(\ z < 0.81) - P(\ z < -0.81) \\\\ ( P( z < 0.81) $ and $P( z < -0.81) $ are closest to 0.79 and 0.21 $ \\ $ respectively from Z table $) \\\\ \Rightarrow (\ -\dfrac{10,150}{\sigma} <z < \dfrac{10,150}{\sigma}\ )= P(\ -0.81 < z < 0.81 \ ) \\\\ \Rightarrow \dfrac{10,150}{\sigma} = 0.81 \hspace{15 mm} ( $ Comparing both sides $ ) \\\\ \Rightarrow \sigma = \dfrac{10,150}{0.81} \approx \textbf{\$12,530.86}